# Low-Rank Quantization-Aware Training for LLMs

### v1

@@ Abstract
Large language models (LLMs) are omnipresent, however their practical deployment is challenging due to their ever increasing computational and memory demands. Quantization is one of the most effective ways to make them more compute and memory efficient. Quantization-aware training (QAT) methods, generally produce the best quantized performance, however it comes at the cost of potentially long training time and excessive memory usage, making it impractical when applying for LLMs. Inspired by parameter-efficient fine-tuning (PEFT) and low-rank adaptation (LoRA) literature, we propose LR-QAT – a lightweight and memory-efficient QAT algorithm for LLMs. LR-QAT employs several components to save memory without sacrificing predictive performance: (a) low-rank auxiliary weights that are aware of the quantization grid; (b) a downcasting operator using fixed-point or double-packed integers and (c) checkpointing. Unlike most related work, our method (i) is inference-efficient, leading to no additional overhead compared to traditional PTQ; (ii) can be seen as a general extended pretraining framework, meaning that the resulting model can still be utilized for any downstream task afterwards; (iii) can be applied across a wide range of quantization settings, such as different choices quantization granularity, activation quantization, and seamlessly combined with many PTQ techniques. We apply LR-QAT to the LLaMA-2/3 and Mistral model families and validate its effectiveness on several downstream tasks. Our method outperforms common post-training quantization (PTQ) approaches and reaches the same model performance as full-model QAT at the fraction of its memory usage. Specifically, we can train a 7B LLM on a single consumer grade GPU with 24GB of memory.
@@ 1 Introduction
We apply LR-QAT to the LLaMA-2/3 and Mistral model families and demonstrate its effectiveness on several general language modeling datasets and zero-shot evaluation on some of the common reasoning downstream tasks. Our method outperforms recent post-training quantization (PTQ) approaches and reaches the same predictive performance as full-model QAT at the fraction of its memory usage. Finally, our method can be applied across a wide range of quantization settings (per-channel or per-block weight quantization, activation quantization, etc.) and can be combined with most of other PTQ techniques.
@@ 2 Background and related work
Figure 1: Left: Visual illustration of our proposed LR-QAT for memory efficient QAT. Right: Memory requirements for various QAT techniques on LLaMA-2 7B (assuming batch size 1 and a sequence length of 1024).
@@
Low-rank adaptation (LoRA) [24] is a parameter efficient fine-tuning (PEFT) method that reduces memory requirements compared to standard training. LoRA freezes the pretrained weights $\bm{W}=\bm{W_{0}}$, and only trains a small set of low-rank trainable parameters, often termed adapters. Given a linear projection $\mathbf{y}=\bm{W}\mathbf{x}$ with $\bm{W}\in\mathbb{R}^{m\times k}$, LoRA computes
$\mathbf{y}=\bm{W}\mathbf{x}+\frac{\alpha}{r}\bm{A}\bm{B}\mathbf{x},$ (2)
@@
Naturally, there have been several works that explored the combination of LoRA and quantization. QLoRA [16] quantizes the pretrained weights to 4 bit using NF4 format and dequantizes them in the forward pass to further reduce fine-tuning memory footprint. QA-LoRA [63] uses INT4 quantization and introduces group-wise operators to enable quantization during inference stage. LoftQ [37] proposed an iterative SVD-based procedure for initializing $\bm{A}$, $\bm{B}$ that yields faster fine-tuning convergence when used together with low-bit quantization. LQ-LoRA [20] further extends initialization technique from LoftQ to mixed precision and data aware cases. Other recent works include [28, 67].
@@
PTQ, while generally fast and simple, suffers from limited performance in low-bit scenarios. Although QAT methods still perform well in low-bit regimes, their high training costs and memory usage make them impractical for LLMs.
LoRA-based methods address memory issues for efficient fine-tuning. However, in most cases they don’t consider efficient inference. The adapters $\bm{A}$, $\bm{B}$ are typically stored in higher precision format such as BF16 and at inference they dequantize the low-bit integer matrix $\bm{W}_{\mathbb{Z}}$ to the same data format, resulting in runtime overhead.
Simply quantizing adapters after training will lead to a different quantization grid compared to $\bm{W}$, and quantizing them using the same quantization grid as $\bm{W}$ will lead to high error. QA-LoRA is the only work we are aware of that attempts to fuse auxiliary LoRA weights back into the frozen $\bm{W}_{\mathbb{Z}}$. Yet, their method is designed to only work with group-wise quantization with high number of groups (a small group size of $32$). In addition to that, QA-LoRA and most of LoRA-based methods combine their proposed technique with the task-specific fine-tuning, whereas we propose LR-QAT as an extended pretraining method.
@@
Table 1: Comparison between existing approaches and the proposed method.
@@ 3 Method
where weights $\bm{W}$ (often termed “shadow weights”) are trainable parameters and the quantization scale $\mathbf{s}$ can be either fixed or also learned. To be able to backpropagate through round-to-nearest operation in (3), it is common to use straight-through estimator (STE, Bengio et al. 3), where it is assumed that $\frac{\partial\!\left\lfloor{t}\right\rceil}{\partial t}=1$. When applied to LLMs, it is straightforward to see that this procedure is very expensive: we have to learn a comparable number of parameters that was used for pretraining, leading to excessive memory usage.
To make this approach more practical we freeze the pretrained weights $\bm{W}$ (denote $\bm{W_{0}}$) and introduce low-rank adapters $\bm{A}\in\mathbb{R}^{m\times r}$, $\bm{B}\in\mathbb{R}^{r\times k}$, $r\ll\min\left\{m,k\right\}$. We have to be careful where exactly those adapters are placed. As discussed in Section 2, after the training is complete, we want $\bm{A}$ and $\bm{B}$ to be seamlessly integrated into a single $b$-bit integer matrix $\bm{W}_{\mathbb{Z}}$ without loss of accuracy to facilitate the efficient inference. To accommodate that, we put the auxiliary matrices inside the rounding operator as follows
@@
where we are using STE assumption for the rounding operation to compute the gradients of the loss w.r.t. $\bm{A}$, $\bm{B}$ and $\mathbf{s}$. We further employ a scaling factor $\alpha/r$ used in LoRA [24] to reduce the need to retune hyperparameters as we vary the rank $r$. After training is complete, (4) can be represented as regular fixed point tensor, $\widehat{\bm{W}}=\mathbf{s}\cdot\bm{W}_{\mathbb{Z}}$, without any extra effort or loss of accuracy and therefore enabling efficient inference without any extra overhead. Note that this is different to most of the literature, such as QLoRA [16], where adapters are placed outside of the quantization function (such as $\mathbf{y}=\widehat{\bm{W}}\mathbf{x}+\bm{A}\bm{B}\mathbf{x}$) and typically stored in higher precision formats such as BF16.
@@
Given that the weight matrix $\bm{W_{0}}$ is frozen, the next natural step to further reduce the memory utilization is to store it in a lower-precision format. One could apply downcasting of $\bm{W_{0}}$ directly in (4), however notice that those weights are divided by the scale $\mathbf{s}$ at every forward pass, which generally has to be stored in a high-precision format to guarantee stable training. To simplify further, we propose the following variant of low-rank QAT:
@@
where $\varphi\left(\cdot\right)$ – a downcasting operator that encapsulates a choice of different numeric formats or other preprocessing computations. In the simplest form, $\varphi\left(\cdot\right)$ would cast the input to one of pre-existing floating-point formats, such as FP16, BF16, FP8 etc.
@@
Another way is to use INT8 storage type, allocate $b$ bits to represent the integer part as before, but utilize the remaining $8-b$ bits for storing the approximate fractional part ($2\leq b\leq 7$). In other words, we represent $\bm{\Phi}_{0}$ using fixed-point numbers. Specifically, assuming the the rest of the computation is performed in BF16, we define the downcasting and the corresponding upcasting operators as follows:
@@
A fixed-point number where $n$ bits are used for the integer part of the value and $m$ bits are used for the fractional part are commonly denoted [50] as Qn.m. For brevity, we will refer to (7) as $\varphi={\text{{\small{Q$b$.$(8-b)$}}}}$. Specifically, in this work we will be mainly focusing on $b\in\left\{3,4\right\}$, which corresponds to Q3.5 and Q4.4, respectively.
@@
Note that both in the original LoRA paper [24] and in the related work like QLoRA [16], there is no need to compute the product $\bm{A}\bm{B}$. Instead, those matrices are multiplied with the activations $\mathbf{x}$ as $\bm{A}\left(\bm{B}\mathbf{x}\right)$. However, we do compute a product $\bm{A}\bm{B}$ in (5), and in the naive implementation of our method, this product together with the results of some intermediate computations (e.g., after rounding and clipping) will be automatically kept in memory for the backward pass, leading to increased memory usage. To prevent this, we employ gradient checkpointing [10] on (5). In other words, we recompute the quantizer function in the backward pass, leading to a slight runtime overhead but avoiding significantly increased memory usage.
@@ 4 Experiments
We assess the effectiveness of LR-QAT by conducting experiments on LLaMA-2 7B/13B [58], LLaMA-3 8B [1] and Mistral-0.1 7B [30]. To compare to relevant literature with weight-activation quantization, we further experiment with LLaMa-1 7B [57]. We first explore the impact of the choice of rank $r$, a downcasting operator $\varphi\left(\cdot\right)$, and the initialization of auxiliary matrices $\bm{A}$, $\bm{B}$. We then compare our method in terms of accuracy to standard full-model QAT, other baselines, and the related work. All detailed hyperparameters of our experiments are in Appendix A.
Quantization
We experiment with both weight-only and weight-activation quantization. The default settings are INT4 / INT3 per-channel (denoted ‘pc’) and group-wise weight quantization with a group size of 128 (denoted ‘g128’). We quantize all linear layers, except the classification head. In weight-activation quantization, defaults are INT4 per-channel weight and per-token activation quantization [14]. Following OmniQuant [52] we quantize all inputs to matmuls with exception of the softmax output and additionally quantize the KV-cache as in LLM-QAT [42].
Evaluation
Datasets and training
We train on a small subset of SlimPajama [54], which is a close open-source replica of the dataset used for pretraining LLaMA models. For weight-only quantization results, we train for $10^{4}$ steps using batch size $32$ and a maximum sequence length of $1024$. For ablation studies and weight-activation quantization results, we use shorter training of $10^{3}$ steps. We select hyperparameters based on the perplexity of a small subset of Wikipedia validation set (512 sequences).
Baselines
We compare with round-to-nearest quantization (RTN), where we set the ranges based on minimizing the $L^{p}$-norms between quantized and unquantized weights and report the best performing configuration. We also use that as initialization for LR-QAT. For weight-only quantization, we compare with GPTQ [18], AWQ [39], OmniQuant [52], and our implementation of PEQA [31], where we use symmetric weight quantization and following the same experimental setup and best RTN initialization, for a fair comparison. For weight-activation quantization, we compare to RTN, SmoothQuant [62] and LLM-QAT [42]. Following [42], we compare to them in several different settings, where the weights, activations and KV cache values are quantized to different levels (denoted as W-A-KV).
4.1 The impact of rank $r$
@@
$\varphi\left(\cdot\right)$ dtype $\bm{A}$, $\bm{B}$ init. W4 @ pc W3 @ pc
@@
4.2 The impact of downcasting operator $\varphi\left(\cdot\right)$ and initialization
@@
Finally, when using the fixed point representation for $\bm{\Phi}_{0}$, we still maintain the same model performance by switching the compute data type [^3] A data type used for activations, gradients, and frozen parameters. from FP32 to BF16, where the latter is what is commonly used for LLMs.
4.3 Comparison with full-model QAT
Table 3: Comparison of the proposed method with the vanilla (full-model) QAT on LLaMA-2 7B with W4 per-channel quantization.
Method GPU memory, GB WikiText-2 perplexity $\downarrow$ Zero-shot avg. accuracy $\uparrow$
Full model QAT 71.9 5.74 69.29
LR-QAT 20.5 5.68 69.43
[…]
4.4 Main results
Table 4: Weight-only quantization results for LLaMA-2/3 and Mistral models. We report WikiText-2 test set perplexity (lower is better) and zero-shot average accuracy (higher is better). Models marked ‘L2’, ‘L3’, and ‘M’ denote LLaMA-2, LLaMA-3, and Mistral, respectively. Numbers marked in bold are SOTA or on par (within $0.05$). §Uses asymmetric weight quantization.
Bitwitdth Method WikiText-2 Perplexity $\downarrow$ Zero-shot avg. accuracy $\uparrow$

### v2

@@ Abstract
Large language models (LLMs) are omnipresent, however their practical deployment is challenging due to their ever increasing computational and memory demands. Quantization is one of the most effective ways to make them more compute and memory efficient. Quantization-aware training (QAT) methods, generally produce the best quantized performance, however it comes at the cost of potentially long training time and excessive memory usage, making it impractical when applying for LLMs. Inspired by parameter-efficient fine-tuning (PEFT) and low-rank adaptation (LoRA) literature, we propose LR-QAT – a lightweight and memory-efficient QAT algorithm for LLMs. LR-QAT employs several components to save memory without sacrificing predictive performance: (a) low-rank auxiliary weights that are aware of the quantization grid; (b) a downcasting operator using fixed-point or double-packed integers and (c) checkpointing. Unlike most related work, our method (i) is inference-efficient, leading to no additional overhead compared to traditional PTQ; (ii) can be seen as a general extended pretraining framework, meaning that the resulting model can still be utilized for any downstream task afterwards; (iii) can be applied across a wide range of quantization settings, such as different choices quantization granularity, activation quantization, and seamlessly combined with many PTQ techniques. We apply LR-QAT to LLaMA-2/3 and Mistral model families and validate its effectiveness on several downstream tasks. Our method outperforms common post-training quantization (PTQ) approaches and reaches the same model performance as full-model QAT at the fraction of its memory usage. Specifically, we can train a 7B LLM on a single consumer grade GPU with 24GB of memory.
@@ 1 Introduction
We apply LR-QAT to LLaMA-2/3 and Mistral model families and demonstrate its effectiveness on several general language modeling datasets and zero-shot evaluation on some of the common reasoning downstream tasks. Our method outperforms recent PTQ approaches and reaches the same predictive performance as full-model QAT at the fraction of its memory usage. Finally, our method can be applied across a wide range of quantization settings, including per-channel or per-block weight quantization, activation quantization, and can be combined with most of other PTQ techniques.
@@ 2 Background and related work
Figure 1: Left: A schematic illustration of our proposed LR-QAT. $\mathbf{x}$ and $\mathbf{y}$ denote the input and the output of the linear layer. Right: Memory requirements for training with various QAT techniques on LLaMA-2 7B (assuming batch size 1, sequence length 1024, $r=32$, and BF16 compute data type).
@@
Low-rank adaptation (LoRA) [25] is a parameter efficient fine-tuning (PEFT) method that reduces memory requirements compared to standard training. LoRA freezes the pretrained weights $\bm{W}=\bm{W_{0}}$, and only trains a small set of low-rank trainable parameters, often termed adapters. Given a linear projection $\mathbf{y}=\bm{W_{0}}\mathbf{x}$ with $\bm{W_{0}}\in\mathbb{R}^{m\times k}$, LoRA computes
$\mathbf{y}=\bm{W_{0}}\mathbf{x}+\frac{\alpha}{r}\bm{A}\bm{B}\mathbf{x},$ (2)
@@
Naturally, there have been several works that explored the combination of LoRA and quantization. QLoRA [16] quantizes the pretrained weights to 4 bit using (a non-uniform) NF4 format and dequantizes them in the forward pass to further reduce fine-tuning memory footprint. QA-LoRA [66] uses INT4 quantization and introduces group-wise operators to enable quantization during inference stage. LoftQ [39] proposed an iterative SVD-based procedure for initializing $\bm{A}$, $\bm{B}$ that yields faster fine-tuning convergence when used together with low-bit quantization. LQ-LoRA [21] further extends initialization technique from LoftQ to mixed precision and data aware cases. Other recent works include [29, 70].
@@
While generally fast and simple, PTQ suffers from limited performance in low-bit scenarios. Although QAT methods still perform well in low-bit regimes, their high training costs and memory usage make them impractical for LLMs.
LoRA-based methods address memory issues for efficient fine-tuning. However, in most cases they don’t consider efficient inference. The adapters $\bm{A}$ and $\bm{B}$ are typically stored in higher precision formats, such as BF16. During inference, the frozen low-bit pretrained weights $\bm{W_{0}}$ are dequantized to match the same data format, resulting in runtime overhead.
Simply quantizing adapters after training will lead to a different quantization grid compared to $\bm{W_{0}}$, and quantizing them using the same quantization grid as $\bm{W_{0}}$ will lead to a high error. QA-LoRA is the only work we are aware of that attempts to fuse auxiliary LoRA weights back into the frozen $\bm{W}_{\mathbb{Z}}$. Yet, their method is designed to only work with group-wise quantization with high number of groups (a small group size of $32$). In addition to that, QA-LoRA and most of LoRA-based methods combine their proposed techniques with the task-specific fine-tuning, whereas we propose LR-QAT as an extended pretraining method.
@@
Table 1: A comparison between existing approaches and the proposed method.
@@ 3 Method
where weights $\bm{W}$ are trainable parameters and the quantization scale $\mathbf{s}$ can be either fixed or also learned. To be able to backpropagate through round-to-nearest operation in (3), it is common to use straight-through estimator (STE, Bengio et al. 3), where it is assumed that $\frac{\partial\!\left\lfloor{t}\right\rceil}{\partial t}=1$. When applied to LLMs, it is straightforward to see that this procedure is very expensive: we have to learn a comparable number of parameters that was used for pretraining, leading to excessive memory usage.
To make this approach more practical we freeze the pretrained weights $\bm{W}$ (denote $\bm{W_{0}}$) and introduce low-rank adapters $\bm{A}\in\mathbb{R}^{m\times r}$, $\bm{B}\in\mathbb{R}^{r\times k}$, $r\ll\min\left\{m,k\right\}$. We have to be careful where exactly those adapters are placed. As discussed in Section 2, after the training is complete, we want $\bm{A}$ and $\bm{B}$ to be seamlessly integrated into a single $b$-bit integer matrix $\bm{W}_{\mathbb{Z}}$ without loss of accuracy to facilitate efficient inference. To accommodate that, we put the auxiliary matrices inside the rounding operator as follows
@@
where we are using STE assumption for the rounding operation to compute the gradients of the loss w.r.t. $\bm{A}$, $\bm{B}$ and $\mathbf{s}$. We further employ a scaling factor $\alpha/r$ used in LoRA [25] to reduce the need to retune hyperparameters as we vary the rank $r$. After training is complete, (4) can be represented as regular fixed point tensor, $\widehat{\bm{W}}=\mathbf{s}\cdot\bm{W}_{\mathbb{Z}}$, without any extra effort or loss of accuracy and therefore enabling efficient inference without any extra overhead. Note that this is different to most of the literature, such as QLoRA [16], where adapters are placed outside of the quantization function (such as $\mathbf{y}=\widehat{\bm{W}}\mathbf{x}+\bm{A}\bm{B}\mathbf{x}$) and are typically stored in higher precision formats such as BF16.
@@
Given that the weight matrix $\bm{W_{0}}$ is frozen, the next natural step to further reduce the memory utilization is to store it in a lower-precision format. One could directly apply downcasting to $\bm{W_{0}}$ in (4). However, it’s important to note that these weights are divided by the scale $\mathbf{s}$ during every forward pass. To ensure stable training, the scale generally needs to be stored in a high-precision format. Therefore, to simplify further, we propose the following variant of low-rank QAT:
@@
where $\varphi\!\left(\cdot\right)$ is a downcasting operator that encapsulates a choice of different numeric formats or other preprocessing computations. In the simplest form, $\varphi\!\left(\cdot\right)$ would cast the input to one of pre-existing floating-point formats, such as FP16, BF16, FP8 etc.
@@
Another way is to use INT8 storage type, allocate $b$ bits to represent the integer part as before, but utilize the remaining $8-b$ bits for storing the approximate fractional part ($2\leq b\leq 7$). In other words, we represent $\bm{\Phi_{0}}$ using fixed-point numbers. Assuming the rest of the computation is performed in BF16, we define the downcasting and the corresponding upcasting operators as follows:
@@
A fixed-point number where $n$ bits are used for the integer part of the value and $m$ bits are used for the fractional part are commonly denoted [52] as Qn.m. For brevity, we will refer to (7) as $\varphi={\text{{\small{Q$b$.$(8-b)$}}}}$. In this work we will be mainly focusing on $b\in\left\{3,4\right\}$, which corresponds to Q3.5 and Q4.4, respectively.
@@
Note that both in the original LoRA paper [25] and in the related work like QLoRA [16], there is no need to compute the product $\bm{A}\bm{B}$. Instead, those matrices are multiplied with the activations $\mathbf{x}$ as $\bm{A}\left(\bm{B}\mathbf{x}\right)$. However, we do compute a product $\bm{A}\bm{B}$ in (5), and in the naïve implementation of our method, this product together with the results of some intermediate computations (e.g., after rounding and clipping) will be automatically kept in memory for the backward pass, leading to increased memory usage. To prevent this, we employ gradient checkpointing [10] on (5). In other words, we recompute the quantizer function in the backward pass, leading to a slight runtime overhead but avoiding significantly increased memory usage.
@@ 4 Experiments
$\varphi\!\left(\cdot\right)$ dtype $\bm{A}$, $\bm{B}$ init. W4pc W3pc
@@
We assess the effectiveness of LR-QAT by conducting experiments on LLaMA-2 7B/13B [61], LLaMA-3 8B [1] and Mistral-0.1 7B [31]. To compare to relevant literature with weight-activation quantization, we further experiment with LLaMa-1 7B [60]. We first explore the impact of the choice of rank $r$, a downcasting operator $\varphi\!\left(\cdot\right)$, and the initialization of auxiliary matrices $\bm{A}$, $\bm{B}$. We then compare our method in terms of accuracy to standard full-model QAT, other baselines, and the related work. All detailed hyperparameters of our experiments are in Appendix B.
Quantization
We experiment with both weight-only and weight-activation quantization. The default settings are INT4 / INT3 per-channel (denoted ‘pc’) and group-wise weight quantization with a group size of 128 (denoted ‘g128’). We quantize all linear layers, except the classification head. In weight-activation quantization, defaults are INT4 per-channel weight and per-token activation quantization [14]. Following OmniQuant [55], we quantize all inputs to matmuls with exception of the softmax output and additionally quantize the KV-cache as in LLM-QAT [44].
Evaluation
Datasets and training
We apply our methods to all linear layers in the attention blocks (both in self-attention and in the feed-forward network). We only train low-rank auxiliary matrices $\bm{A}$, $\bm{B}$ and the quantization parameters $\mathbf{s}$. We train on a small subset of SlimPajama [57], which is an open-source dataset similar to the original one used for pretraining LLaMA models. In all experiments we train using batch size $32$ and a maximum sequence length of $1024$. For weight-only and weight-activation quantization results, we train for $10^{4}$ steps. For ablation studies, we use shorter training of $10^{3}$ steps. We select hyperparameters based on the perplexity of a small subset of Wikipedia validation set (512 sequences).
Baselines
We compare with round-to-nearest quantization (RTN), where we set the ranges based on minimizing the $L^{p}$-norms between quantized and unquantized weights and report the best performing configuration. We also use that as initialization for LR-QAT. For weight-only quantization, we compare with GPTQ [18], AWQ [41], OmniQuant [55], and our implementation of PEQA [32], where we use symmetric weight quantization and following the same experimental setup and best RTN initialization, for a fair comparison. Similar to our method, we also apply PEQA only to linear layers in the attention blocks and keep token embeddings, final classification head and LayerNorm parameters frozen.
For weight-activation quantization, we compare our method with RTN, SmoothQuant [65], LLM-QAT [44], Outlier Suppression$+$ [63], OmniQuant [55], and our implementation of PEQA [32]. Following [44], we compare to them in several different settings, where the weights, activations and KV cache values are quantized to different bitwidths (denoted as W-A-KV).
[…]
Method GPU memory, GB WikiText-2 perplexity $\downarrow$ Avg. zero-shot accuracy $\uparrow$
Full-model QAT 71.9 5.74 69.29
LR-QAT 20.5 5.68 69.43
4.1 The impact of rank $r$
@@
4.2 The choice of the downcasting operator $\varphi\!\left(\cdot\right)$ and $\bm{A}$, $\bm{B}$ initialization
@@
Finally, when using the fixed point representation for $\bm{\Phi_{0}}$, we still maintain the same model performance by switching the compute data type [^3] A data type used for activations, gradients, and frozen parameters. from FP32 to BF16, where the latter is what is commonly used for LLMs. Going forward, we use $\varphi={\text{{\small{Q$b$.$(8-b)$}}}}$ with ‘LoRA’ initialization and BF16 compute data type.
Table 4: Weight-only quantization results for LLaMA-2/3 and Mistral models. We report WikiText-2 test set perplexity (lower is better) and average zero-shot accuracy (higher is better). Models marked ‘L2’/‘L3’, and ‘M’ denote LLaMA-2/3 and Mistral, respectively. Numbers marked in bold are SOTA or on par (within $0.05$). §Uses asymmetric weight quantization.
# Bits Method WikiText-2 Perplexity $\downarrow$ Avg. zero-shot accuracy $\uparrow$
@@
4.3 Comparison with full-model QAT
[…]
4.4 Main results